what is homogeneous linear equation &nonhomogeneous linear equation?what is consistence&incossistence?


1
Expert's answer
2021-07-08T16:39:55-0400

A homogeneous system of linear equations is of the form below for two variables

"ax_1+bx_2=0\\
cx_1+dx_2=0\\"

i.e the values of the variables when substituted into any of the equation give zero.

However, non-homogeneous system of linear equations is of the form

"ax_1+bx_2=d_1\\
cx_1+dx_2=d_2"

where both "d_1,d_2" are non-zero.

For three variables, homogeneous system of linear equations is of the form

"ax_1+bx_2+cx_3=0\\
dx_1+ex_2+fx_3=0\\
gx_1+hx_2+jx_3=0\\"

while a form of non-homogeneous system of linear equations is

"ax_1+bx_2+cx_3=d_1\\

dx_1+ex_2+fx_3=d_2\\

gx_1+hx_2+jx_3=d_3\\"

If any of the above systems of equations has a solution then such a system of equations is said to be consistent. If the system of equations does not have solution, then the system is inconsistent.

Consider the equations below

"x+y=5\\
-2x-2y=6\\"

x and y cannot cancel out due to the reoccurrence of the first equation in the second equation. i.e.

"-2(x+y)=6"

So it's inconsistent.

However, the equations below are consistent.

"x+y=5\\
-x+y=1\\"

Because x=2 and y=3.